Gordon's age is 7 more than 5 times Hank's age. If Hank is 4 years old, how old is Gordon?
Found 2 solutions by ikleyn, amalm06:
Answer by ikleyn(52781)

.
Gordon's age = 5*4+7 years.


Answer by amalm06(224)

5H+7=G

H=4 ---> G=(5)(4)+7=27 (Answer)